# End-to-End Speech-to-Text Translation: A Survey

### v1

@@ Abstract
Speech-to-text translation pertains to the task of converting speech signals in a language to text in another language. It finds its application in various domains, such as hands-free communication, dictation, video lecture transcription, and translation, to name a few. Automatic Speech Recognition (ASR), as well as Machine Translation(MT) models, play crucial roles in traditional ST translation, enabling the conversion of spoken language in its original form to written text and facilitating seamless cross-lingual communication. ASR recognizes spoken words, while MT translates the transcribed text into the target language. Such disintegrated models suffer from cascaded error propagation and high resource and training costs. As a result, researchers have been exploring end-to-end (E2E) models for ST translation. However, to our knowledge, there is no comprehensive review of existing works on E2E ST. The present survey, therefore, discusses the work in this direction. Our attempt has been to provide a comprehensive review of models employed, metrics, and datasets used for ST tasks, providing challenges and future research direction with new insights. We believe this review will be helpful to researchers working on various applications of ST models.
@@ 1 Introduction
The traditional ST translation methods involve a cascade approach– First, applying automatic speech recognition (ASR) on the given speech and then performing machine translation (MT) on the transcription produced by ASR. Such a cascade model is prone to several issues, such as (a) error in the ASR model can propagate to the MT model, (b) higher cost and training time, and (d) resources required for training. To mitigate such issues, various researchers propose using end-to-end (E2E) models for ST task (Bérard et al., 2016, 2018; Anastasopoulos et al., 2016; Bentivogli et al., 2021; Gangi et al., 2019). E2E models usually employ a single neural network that can be trained end-to-end. Because of simpler training, lower memory footprint, and cost, E2E model development has gained significant momentum in the research community. However, a systematic and comprehensive review of ST works is missing. Concurrently to our work, we find that a review paper (Xu et al., 2023) on ST was published recently. The aforementioned review categorizes existing works mainly based on modeling, data, and application issues. They do not cover the data sets available for the ST tasks nor provide any insights into the cascade vs. E2E model performances. Also, the future open problems provided by them are limited. On the other hand, our work comprehensively reviews the existing models for ST tasks and datasets from a completely different perspective and critically analyzes the existing works; thereafter, we identify several challenges and future research directions. Thus, our work may be deemed complimentary to (Xu et al., 2023).
The following review is structured as follows: In Section 2, we establish the foundation of the ST task through a formal definition, and we subsequently delve into the various metrics and loss functions adopted by different researchers. Moving on to Section 3, we present comprehensive descriptive statistics pertaining to benchmark datasets that have been employed by researchers within this field. Section 4 is dedicated to exploring the different toolkits accessible for addressing E2E ST tasks. In Section 5, we delve into the diverse strategies and approaches employed to tackle the ST problem. Initially, we categorize these approaches based on the frameworks utilized and the characteristics of the data involved. Under the framework categorization in Section 5.1, we provide explanations for both Sequence-to-sequence (seq2seq) frameworks (5.1.1) and Modality-bridging frameworks (5.1.2). Additionally, the seq2seq frameworks are further subcategorized, taking into account scenarios involving ST with MT, ST with ASR, and ST with both MT and ASR. Furthermore, the nature of the data is dissected into several distinctive settings in Section 5.2, including Low-resource (5.2.1), Code-mixed (5.2.2), Streaming (5.2.3), Unsupervised (5.2.4), and Multilingual (5.2.5). Finally, in Section 6, we explore the prospects for future research and address open problems within the field.
@@ 2 Background
This section describes the ST task formally and presents the metrics used to evaluate the performance of the ST model.
@@ 2.1 Task Definition
ST task can be defined as translating the given input speech $U$ in one language to translated text $V$ in another language with the transcription text $X$. The definition, in formal terms, is as follows: Given a dataset $D=\{({\bf u}^{i},{\bf x}^{i},{\bf v}^{i})|i=1,2,\ldots,n\}$ of pairs of input speech features ${\bf u}=(u_{1},u_{2},\ldots,u_{T_{u}})$ in a language and output text tokens ${\bf v}=(v_{1},v_{2},\ldots,v_{T_{v}})$ in a different language, the objective of the ST task is to minimize the conditional probability given below:
@@
In the above equation, $T_{u}$, $T_{v}$, and $\theta$ are the lengths of input features, the number of output tokens, and the model parameter, respectively. There are $n$ parallel sentences in our corpus, and the model is optimized for negative log-likelihood over these sentences as
@@
The above optimization is usually solved using an encoder-decoder with an attention approach. Essentially, an encoder maps speech input to a hidden state representation $h$ followed by a decoder which takes the previously generated text tokens $v_{<t}$, encoder hidden state $h$ and attention vector ${\bf\alpha}$ (Vaswani et al., 2017). Offline ST translation can look at the whole speech before producing output text tokens, whereas online ST can start translation as soon as it sees a few seconds of the speech signal.
2.2 Evaluation Metrics
This section discusses various metrics used to evaluate the E2E ST models. The metrics of E2E ST models are categorized into 2 major performance measures which are quality and latency. The quality of the E2E ST models is the measure of how close the ST translation is to the target sentence or human-translated sentence. The latency of the E2E ST models is the time taken by the model to translate source speech into the target text.
Quality-based metrics: The quality-based metrics measure how close the translation is to the target sentence, which usually lies in [0, 1]. Most of the existing literature evaluates these scores on detokenized output. Standard translation metrics for evaluating ST task performance are presented below:
@@ 2.2 Evaluation Metrics
    BLEU Bi-lingual Evaluation Understudy (BLEU) Papineni et al. (2002) is the score of weighted average length of the translated sentence that matched the target sentence. The expression for evaluating BLEU score is given by:
    $BLEU=BP.exp(\sum_{n=1}^{N}w_{n}\log P_{n})$ (3)
    where the $BP$, $w_{n}$, and $P_{n}$ are brevity penalty, positive weights that sums to 1, and n-gram precisions of length upto $N$, respectively.
  • 2.
    TER Translation Edit Rate (TER) Snover et al. (2006) is the measure of how much a translated sentence has to be modified in order to match the target sentence. The expression for TER is given by:
    $TER=M/R$ (4)
    where $M$ and $R$ are the modifications required in the translated sentence and the average number of reference words, respectively.
  • 3.
    METEOR Metric for Evaluation of Translation with Explicit word Ordering (METEOR) Banerjee and Lavie (2005) is the automatic metric that matches the unigrams in the translated text to the referenced text. The expression for METEOR Score is given by:
    $M\_Score=F_{M}*(1-Penalty_{M}),$ (5)
    where $F_{M}=\frac{10PR}{R+9P},$ and $Penalty_{M}=0.5*\left(\frac{C}{U}\right)^{3}.$
    The $F_{M}$ and $Penalty_{M}$ are the F-Score and the penalty of the METEOR score. $P$, $R$, $C$, and $U$ denote the precision, recall, number of chunks, and number of unigrams matched in the translated text, respectively.
  • 4.
    BERTScore The BERTScore Zhang et al. (2019) is an automatic evaluation metric that scores the similarity between the translated text to the referenced text. It takes into account the Recall ($R$), Precision ($P$), and Fscore ($F$), which are expressed as:
    $R=\frac{1}{|v|}\sum_{v_{i}\in v}\max_{\hat{v_{j}}\in\hat{v}}v_{i}^{T}\hat{v_{j}},P=\frac{1}{|\hat{v}|}\sum_{\hat{v_{j}}\in\hat{v}}\max_{v_{i}\in v}v_{i}^{T}\hat{v_{j}},F=2.\frac{R.P}{R+P}$ (6)
    where $v_{i}$ and $\hat{v_{j}}$ are tokens of the sentences in the translated text and the ground truth text.
There are few other evaluation metrics such as TRANSTAC Schlenoff et al. (2009), NIST-Score Doddington (2002) and CHRF Popović (2015) which are less frequently used.
Latency-based metrics: For streaming ST task, researchers report a metric for measuring latency, which is defined as the delay incurred in starting to produce the translation. Let ${\bf u},{\bf v}$ and $\hat{{\bf v}}$ denote the input speech sequence, ground truth text sequence, and system-generated hypothesis sequence, respectively. In the streaming ST task, the model can only produce output with partial input. Suppose ${\bf u}_{1:t}=(u_{1},\ldots,u_{t}),t<T_{u}$ has been read when generating $v_{s}$, the delay in $v_{s}$ is defined as (Ma et al., 2020)
$d_{s}=\sum_{k=1}^{t}T_{k}$ (7)
where $T_{k}$ is the duration of the speech frame $u_{k}$. The latency metrics are evaluated using a method that analyzes a sequence of time delays $[d_{1},\ldots,d_{T_{v}}]$.
  • 1.
    Average Proportion (AP) calculates the mean fraction of the source input that is read during the target prediction generating process.
    $AP=\frac{1}{T_{y}\sum_{k=1}^{T_{u}}T_{k}}\sum_{s=1}^{T_{v}}d_{s}$ (8)
@@
    Where $\tau(T_{u})=\min\{s\mid d_{s}=\sum_{k=1}^{T_{u}}T_{k}\}$ and $\hat{d_{s}}$ are the delays of an ideal policy defined as (Ma et al., 2020)
    $\hat{d_{s}}=(s-1)\sum_{k=1}^{T_{u}}\frac{T_{k}}{\hat{T_{v}}}$ (10)
@@ 2.3 Loss Functions
This section discusses the various loss functions used for the E2E ST models. Let $D=(u,x,v)$ be a tuple where $u,x$, and $v$ are the speech, the transcription text, and the translation text, respectively. The following are the various loss functions that are used to optimize the performance of the E2E ST models:
@@
    ST Loss (Ouyang et al., 2022) is defined as the negative log-likelihood of the translation text given the source speech as follows
    $L_{ST}=-\sum_{(u,v)\in D}\log p(v|u)$ (13)
@@
    MT Loss (Ouyang et al., 2022) is defined as the negative log-likelihood of the translation text given the source transcript as follows
    $L_{MT}=-\sum_{(x,v)\in D}\log p(v|x)$ (14)
  • 3.
    ASR Loss (Ouyang et al., 2022) is defined as the negative log-likelihood of the transcription text given the source speech as follows
    $L_{MT}=-\sum_{(u,x)\in D}\log p(x|u)$ (15)
  • 4.
@@
  • 5.
    Cross-Modal Adaptation Loss (Liu et al., 2020b) is defined as the sum of all the Mean Squared Errors of the speech and the transcription texts.
@@
  • 7.
  • 8.
    Contrastive Loss (Ye et al., 2022a) It is computed between the the speech and the transcription text bringing them closer, and pushing the unrelated pairs farther.

### v2

@@ Abstract
Speech-to-Text (ST) translation pertains to the task of converting speech signals in one language to text in another language. It finds its application in various domains, such as hands-free communication, dictation, video lecture transcription, and translation, to name a few. Automatic Speech Recognition (ASR), as well as Machine Translation(MT) models, play crucial roles in traditional ST translation, enabling the conversion of spoken language in its original form to written text and facilitating seamless cross-lingual communication. ASR recognizes spoken words, while MT translates the transcribed text into the target language. Such integrated models suffer from cascaded error propagation and high resource and training costs. As a result, researchers have been exploring end-to-end (E2E) models for ST translation. However, to our knowledge, there is no comprehensive review of existing works on E2E ST. The present survey, therefore, discusses the works in this direction. We have attempted to provide a comprehensive review of models employed, metrics, and datasets used for ST tasks, providing challenges and future research direction with new insights. We believe this review will be helpful to researchers working on various applications of ST models.
@@ 1 Introduction
Figure 1: History of E2E ST models. Blue color models correspond to streaming models discussed in §7.1.2. Note that here we have listed only a few selected representative models.
Historically, the ST problem has been solved by pipelining ASR and MT models together where ASR models take speech in a source language as input and generate the transcript. Whereas MT models translate the transcript into the target language. Such a cascade model suffers from problems like error propagation, higher training, and inference latency. Therefore, the current trend in developing the ST model is toward the E2E system which is defined as
Definition 1
A unified E2E ST model is implemented, facilitating combined training and recognition processes aimed at consistently reducing the anticipated error rate, thereby bypassing the need for independently acquired sources of knowledge.
Therefore, the main goal of the E2E ST model is to achieve a reduced error rate, with secondary objectives potentially including decreased training/inference duration and memory usage.
There has been a lot of work building E2E ST models (as shown in fig. 1), datasets, and metrics in recent years. However, a systematic and comprehensive review of E2E ST works is missing. The authors found that a review paper (Xu et al., 2023b) on ST was published recently. The review mentioned above categorizes existing works mainly based on modeling, data, and application issues. They do not cover the data sets available for the ST tasks nor provide any insights into the cascade vs. E2E model performances. Also, the future open problems provided by them are limited. On the other hand, our work comprehensively reviews the existing models for ST tasks, evaluation methods, metrics, and datasets from a completely different perspective and critically analyzes the existing works; after that, we identify several challenges and future research directions. Thus, our work may be deemed complimentary to (Xu et al., 2023b).
Figure 2: Organization of the survey paper
The following review is structured following the taxonomy in fig. 2. In §2, we establish the foundation of the ST task through a formal definition, and we subsequently delve into the various metrics and loss functions adopted by different researchers in §3. A comparative discussion between cascade and end-to-end models is presented in §4. Training of E2E ST models suffers from data issues and how to combat them is elaborated in §5. Speech and Text segmentation and representation is an important task in ST model development discussed in §6. In §7, we delve into the strategies employed to tackle the ST problem. We categorize these approaches based on the frameworks utilized and the characteristics of the data involved. Data and toolkits required for ST modeling are discussed in §9. Finally, in §10, we explore the prospects for future research and open problems within the field.
@@ 2 Background
This section describes the ST task formally and presents the loss functions and evaluation metrics commonly employed to optimize ST models.
@@ 2.1 Task Definition
ST task can be defined as translating the given input speech $U$ in one language to translated text $V$ in another language with the transcription text $X$ (optionally). Formally, it is defined as follows: Given a dataset $D=\{({\bf u}^{i},{\bf x}^{i},{\bf v}^{i})|i=1,2,\ldots,n\}$ of pairs of input speech features ${\bf u}=(u_{1},u_{2},\ldots,u_{T_{u}})$ in a language and output text tokens ${\bf v}=(v_{1},v_{2},\ldots,v_{T_{v}})$ in a different language, the objective of the ST task is to minimize the conditional probability given below:
@@
In the above equation, $T_{u}$, $T_{v}$, and $\theta$ are the lengths of input features, the number of output tokens, and the model parameter, respectively. Note that the problem formulation given in (1) is for Autoregressive (AR) models  [^1] Non-autoregressive (NAR) models are an alternative modeling approach that has been proposed in the past few years for the ST task. Only a sparse number of works exist in the literature. We discuss NAR briefly in §7.1. Usually, it is assumed that there are $n$ parallel speech-text pairs in our corpus, and the model is optimized for negative log-likelihood over these pairs as
@@
The above optimization is usually solved using an encoder-decoder with an attention approach. Essentially, an encoder maps speech input to a hidden state representation $h$ followed by a decoder which takes the previously generated text tokens $v_{<t}$, encoder hidden state $h$ and attention vector ${\bf\alpha}$ (Vaswani et al., 2017). Offline ST translation can look at the whole speech before producing output text tokens, whereas streaming ST can start translation of partial speech signal.
3 Evaluation Metrics
This section discusses various metrics used to evaluate the E2E ST models. The metrics to evaluate E2E ST models are categorized into two types: quality and latency. The quality of the E2E ST models is the measure of how close the ST translation is to the target sentence. The latency is the time elapsed between the pronunciation of a word and the generation of its textual translation.
3.1 Quality-based metrics
The quality-based metrics measure how close the translation is to the target sentence. Most of the existing literature evaluates these scores on detokenized output which is the string formed by combining the tokens. Standard metrics for evaluating ST task performance are commonly used MT evaluation metrics such as Bi-lingual Evaluation Understudy (BLEU) (Papineni et al., 2002), Translation Error Rate (TER) (Snover et al., 2006) via sacreBLEU, Metric for Evaluation of Translation with Explicit word Ordering (METEOR) (Banerjee and Lavie, 2005), and CHaRacter-level F-score (CHRF), and CHRF++ (Popović, 2015). Recently BERTScore has shown promising results on comparing with human evaluations. The BERTScore (Zhang et al., 2019) is an automatic evaluation metric that scores the similarity between the translated text and the referenced text. It takes into account the Recall, Precision, and Fscore. There are a few other evaluation metrics such as TRANSTAC (Schlenoff et al., 2009) and which are less frequently reported.
3.2 Latency-based metrics
For streaming ST tasks, researchers report a metric for measuring latency, which is defined as the delay incurred in starting to produce the translation. Let ${\bf u},{\bf v}$ and $\hat{{\bf v}}$ denote the input speech sequence, ground truth text sequence, and system-generated hypothesis sequence, respectively. In the streaming ST task, models produce output with partial input. Suppose ${\bf u}_{1:t}=\{(u_{1},\ldots,u_{t}),t<T_{u}\}$ has been read when generating $v_{s}$, the delay in $v_{s}$ is defined as (Ma et al., 2020a)
$d_{s}=\sum_{k=1}^{t}T_{k}$ (3)
where $T_{k}$ is the duration of the speech frame $u_{k}$. The latency metrics are calculated using a method that analyzes a sequence of time delays $[d_{1},\ldots,d_{T_{v}}]$.
@@ 3.2 Latency-based metrics
    Average Proportion (AP) (Cho and Esipova, 2016a) calculates the mean fraction of the source input that is read during the target prediction generating process.
    $AP=\frac{1}{T_{v}\sum_{k=1}^{T_{u}}T_{k}}\sum_{s=1}^{T_{v}}d_{s}$ (4)
@@
    Where $\tau(T_{u})=\min\{s\mid d_{s}=\sum_{k=1}^{T_{u}}T_{k}\}$ and $\hat{d_{s}}$ are the delays of an ideal policy defined as (Ma et al., 2020a)
    $\hat{d_{s}}=(s-1)\sum_{k=1}^{T_{u}}\frac{T_{k}}{T_{v}}$ (6)
@@
  • 4.
    Length-Adaptive Average Lagging (LAAL) One issue with AL metric for simultaneous translation is that though it can handle the under-generation [^2] Under/Over-generation problem refers to the length of the generated text compared to the reference translation text. problem, it is unable to handle over-generation and produces biased score. To alleviate this issue, (Papi et al., 2022a) propose LAAL which modifies (6) as
    $\hat{d_{s}}=(s-1)\sum_{k=1}^{T_{u}}\frac{T_{k}}{\max\{T_{v},\hat{T_{v}}\}}$ (9)
    Essentially, it divides (6) by the maximum length of the reference and predicted text. As such, it can handle both over and under-generation problems.
  • 5.
    Average Token Delay (ATD) AL metric does not take into account the length of the partial translation output, i.e., it does not consider the latency caused by longer outputs. To remedy this issue, ATD (Kano et al., 2023), defined below, has been proposed recently.
    $ATD=\frac{1}{T_{v}}\sum_{s=1}^{T_{v}}(T(v_{s})-T(u_{a(s)}))$ (10)
    where
    $a(s)$ $=\min(s-f(s),d_{s})$ (11)
    $f(s)$ $=(s-1)-a(s-1)$ (12)
    $T(\cdot)$ in (10) represents the ending time of each input or output token. The token is a sub-segment in speech, a character, or a word in text. $a(s)$ represents the index of the input token corresponding to $v_{s}$ in the time difference calculation and $a(0)=0$. $f(s)$ in (11) represents how much longer the duration of the previous translation prefix is than that of the previous input prefix.
@@ 3.3 Loss Functions
Let $D=(u,x,v)$ be a tuple where $u,x$, and $v$ are the speech, the transcription text, and the translation text, respectively. The following are the various loss functions that are used to optimize the performance of the E2E ST models:
@@
  • 3.
    Cross-Modal Adaptation Loss (Liu et al., 2020d) is defined as the sum of all the Mean Squared Errors of the speech and the transcription texts.
@@
  • 5.
    Contrastive Loss (Ye et al., 2022a) is computed between the speech and the transcription text bringing them closer, and pushing the unrelated pairs farther.